A distillation column is fed with a stream of ethanol, propanol and butanol

flowing at a total rate of 1350 kg h−1 and containing 20 wt% ethanol. The top product consists of 80% of the ethanol and 50% of the propanol contained in the feed. All of the 500 kg h−1 of butanol in the feed ends up in the bottom product. Calculate the bottom product component mass flow rates.



1
Expert's answer
2021-01-18T01:15:01-0500

Butanol 500 kg/h

Ethanol + propanol = 1350 - 500 = 850

Ethanol = 850 x 0.2 = 170 kg/h

Propanol = (850 - 170) x 0.5 = 340 kg/h